OpenStudy (anonymous):

Laplace method of differential equation when initial condition is not zero? y''+2y'+3y=0 y(3)=2 y'(3)=37 just out of curiosity if anyone could show me how to solve this using laplace

OpenStudy (anonymous):

In this situation, you must define a new function: \[v(t)=y(t+t_0)~~\Rightarrow~~\begin{cases}v(t)=y(t+t_0)\\v'(t)=y'(t+t_0)\\v''(t)=y''(t+t_0)\end{cases}\] So if you were given the general equation \(ay''+by'+c=g(t)\) (with initial conditions \(y(t_0)=y_0\) and \(y'(t_0)={y_0}'\)), then the transformation gives you \(av''+bv'+c=g(t+t_0)\) (with initial conditions \(v(0)=y(0+t_0)=y_0\) and \(v'(0)=y'(0+t_0)={y_0}'\)). So, \[y''+2y'+3y=0~~\iff~~v''+2v'+3v'=0\] where \(g(t)=g(t+t_0)=0\). \[\begin{align*}v''+2v'+3v&=0\\ \mathcal{L}\left\{v''+2v'+3v\right\}&=0\\ \left(s^2\mathcal{L}\left\{v\right\}-sv(0)-v'(0)\right)+2\left(s\mathcal{L}\left\{v\right\}-v(0)\right)+3\mathcal{L}\left\{v\right\}&=0\\ \left(s^2+2s+3\right)\mathcal{L}\left\{v\right\}-sv(0)-v'(0)-2v(0)&=0\\ \left(s^2+2s+3\right)\mathcal{L}\left\{v\right\}-2s-37-4&=0\\ \mathcal{L}\left\{v\right\}&=\frac{2s+41}{s^2+2s+3}\end{align*}\] After you find \(v(t)\), keep in mind that you're looking for a solution for \(y(t)\), so remember to back-substitute: \[v(t)=y(t+t_0))~~\iff~~v(t-t_0)=y(t)\]

OpenStudy (anonymous):

could you elaborate on the substitution of sv(0) and -v'(0) where s(v(0)) becomes +2s and v'(0) to +3

OpenStudy (anonymous):

\[sv(0)=sy(0+t_0)=sy(t_0)=sy(3)=2s\] \[v'(0)=y'(0+t_0)=y'(t_0)=y'(3)=37\] I was just plugging in the given conditions.

OpenStudy (anonymous):

Or do you mean how I got the term \((\cdots)\mathcal{L}\{y\}\)?

OpenStudy (anonymous):

ohh... that made sense, I miss read it. thanks, I'm going to have to look more into the substitution you made